Please help me do 2 questions! I can't do that! Thank you.r 6. 0/1 points | Previous Answers TanApMath7 4.2.021 My Notes Ask Your Teacher Luis has丰190,000 in his retirement account at
Answer #1

Question 1

use compound interest formula and annuity formula

FV=PV\left(1+\frac{r}{k}\right)^{nk}+PMT\left[\frac{\left(1+\frac{r}{k}\right)^{nk}-1}{\frac{r}{k}}\right]

PV=190000

PMT=2000

k=4 for quarterly compound

n=30 years

r=4.5% = 0.045

FV = 190000 ( + 0,045、304 (1+ 0.045) 3041 +2000

FV 190000 (1+0.011252000 0.01125

FV 190000 (1.01125)122000 мною(1.01125)», 0.01125

3.828-1 FV190000 3.828+2000 0.01125

2.828 FV190000 3.828+2000

FV- 190000 3.828 + 2000 251.4186

FV-727407.46787+ 502837.39684

Ft-= 1230244.8............after 30 years

.

.

.

.

.

.

.

.

.

Question 2

L=PMT\left[\frac{1-\left(1+\frac{r}{k}\right)^{-t}}{\frac{r}{k}}\right]

L= loan amount

PMT=440

k=12 for monthly compound

t =36 months  

r=8% = 0.08

.

L=440\left[\frac{1-\left(1+\frac{0.08}{12}\right)^{-36}}{\frac{0.08}{12}}\right]

L=440\left[\frac{1-\left(1+0.00667\right)^{-36}}{0.00667}\right]

L = 440 | 1-(1.00667)-36 0.00667 し 440

L=440\left[\frac{1-0.78725}{0.00667}\right]

L=440\cdot 31.89505

{\color{Red} L=14041.194}............loan amount

.

here

DP=down payment = 6000

so cash price is

price =DP+L

price-600014041.19

price = 20041.19

{\color{Red} price \approx 20041.2}